A Foucault pendulum is used to demonstrate the rotation of the Earth. They can usually be found in science museums. A certain pendulum has a 120-kg bob swinging on a 16.2-m-long wire with a maximum amplitude of 6.3°.

A. Sketch the pendulum at its maximum displacement and label all known quantities.

B. What is the difference in gravitational potential energy between the maximum displacement and the equilibrium point?

C. What is the maximum speed of the bob? D. Where in the oscillation is the bob when it has its maximum speed?

E. What is the period and frequency of the oscillation?
Answer #1

given
L = 16.2 m
m = 120 kg
theta = 6.3 degrees

A)

mgcoso ho --- ------ mgsino

B) the difference in gravitational potential energy between the maximum displacement and the equilibrium point,

delta_U = m*g*h

= m*g*L*(1 - cos(theta) )

= 120*9.8*(1 - cos(6.3))

= 7.10 J <<<<<<<<<<------------------Answer

C) KE_max = delta_U

(1/2)*m*v_max^2 = delta_U

v_max = sqrt(2*delta_U/m)

= sqrt(2*7.10/120)

= 0.344 m/s <<<<<<<<<<------------------Answer

D) The bob will have maximum speed at lowest position.

E) Time period, T = 2*pi*sqrt(L/g)

= 2*pi*sqrt(16.2/9.8)

= 8.08 s <<<<<<<<<<------------------Answer

frequency, f = 1/T

= 1/8.08

= 0.124 Hz <<<<<<<<<<------------------Answer